OpenStudy (maeganr):

Homework Help ASAP. 5 Problems!

OpenStudy (maeganr):

"Solve each system using elimination. Tell whether the system has one solution, infinitely solutions, or no solution." #19. x-3y=-7 2x=6y-14 #20. 3x-5y=-2 x+3y=4 #21. x+2y=6 2x-4y=-12 #22. 5x+y=15 3y=-15x+6 #24. 3x-y=-2 -2x+2y=8 HELP. THIS IS DUE TOMORROW.!.

OpenStudy (anonymous):

3x-5y=-2 <-- Multiply by 3 x+3y=4 <---Multiply by 5 9x-15y=-6 5x+15y=20 ------------ 14x =14 Then, 14x=14, x=1 Do you see how I got that? :o)

OpenStudy (texaschic101):

19. x - 3y = -7 2x = 6y - 14 -->2x - 6y = - 14 --------------- now your problem reads.... x - 3y = -7 -->(-2)x - 3y = -7 2x - 6y = -14 --------------- -2x + 6y = 14 (result of multiplying by -2) 2x - 6y = -14 ------------- this has no solution 20. 3x - 5y = -2 x + 3y = 4 --->(-3)x + 3y = 4 ---------------- 3x - 5y = -2 -3x - 9y = - 12 (result of multiplying by -3) ---------------add -14y = -14 (divide by -14 to get y by itself) -14y/-14 = -14/-14 y = 1 now sub 14 in for y in either of original equations... x + 3y = 4 x + 3(1) = 4 x + 3 = 4 x = 4 - 3 x = 1 check... x + 3y = 4 1 + 3(1) = 4 4 = 4 (correct) one solution (1,1) ---> x = 1 and y = 1 now you see how it is done....see what you can do

OpenStudy (texaschic101):

I was wrong about # 19....it has infinite solutions....sorry about that.